OpenStudy (anonymous):

simplify 13x-8xy+12xyz+6(3x-5xy-7xyz

OpenStudy (anonymous):

13x - 8xy + 12xyz + 6(3x - 5xy - 7xyz) = 12x - 8xy + 12xyz + 18x - 30xy - 42xyz = 30x - 38xy - 30xyz = 2x(15 - 19y - 15yz)

OpenStudy (anonymous):

31x_38xy-30xyz

OpenStudy (anonymous):

oh yeah, 31x. not 30.

OpenStudy (anonymous):

jamesm u wrote 12x in the 2nd step instead pf 13x

OpenStudy (anonymous):

31x-38xy-30xyz